# thanks: Contributed Equally

### v1

@@ Acknowledgements
We acknowledge that challenging academic experiences can sometimes lead to unexpected collaborations and new perspectives. In this spirit, SH and SS recognize the broader research journey—including an earlier period associated with Prof. Heinz Koeppl (TU Darmstadt)–that ultimately enabled the completion of this work. SS acknowledges Archishman Raju for insightful discussions and critical suggestions, without which the project would have been incomplete. SS also acknowledges Yuuki Matsushita (NCBS–TIFR) and Kunihiko Kaneko (University of Tokyo & Niels Bohr Institute) for valuable discussions. We acknowledge ChatGPT for grammatical corrections and improvements.
@@ Appendix A Treating Noise Expectation
$=\sum_{n=0}^{\infty}\frac{\left(g\alpha\right)^{2n}}{(2n)!}\sum_{p\in\text{permut.}}\prod_{i,j\in p}\left\{2\int_{0}^{t}\int_{-v}^{v}C_{\eta}(u)\exp\left(v-2t\right)\mathrm{d}u\mathrm{d}v\right\}$
$=\sum_{n=0}^{\infty}\frac{\left(g\alpha\right)^{2n}}{n!}\underbrace{\left\{\int_{0}^{t}\exp\left(v-2t\right)\left[\int_{0}^{v}C_{\eta}(u)\mathrm{d}u\right]\mathrm{d}v\right\}^{n}}_{\tilde{C}^{n}(t)}$

### v2

@@ Acknowledgements
We acknowledge that challenging academic experiences can sometimes lead to unexpected collaborations and new perspectives. In this spirit, SH and SS recognize the broader research journey–including an earlier period associated with Prof. Heinz Koeppl (TU Darmstadt)–that ultimately enabled the completion of this work. SS acknowledges Archishman Raju for insightful discussions and critical suggestions, without which the project would have been incomplete. SS also acknowledges Yuuki Matsushita and Kunihiko Kaneko for valuable discussions. We acknowledge ChatGPT for grammatical corrections and improvements.
Data Availability
The data that support the findings of this study are available within the article.
@@ Appendix A Treating Noise Expectation
$=\sum_{n=0}^{\infty}\frac{\left(g\alpha\right)^{2n}}{(2n)!}\sum_{p\in\text{permut.}}\prod_{i,j\in p}\left\{2\int_{0}^{t}\int_{-\tau}^{\tau}C_{\eta}(s)\exp\left(\tau-2t\right)\mathrm{d}s\mathrm{d}\tau\right\}$
$=\sum_{n=0}^{\infty}\frac{\left(g\alpha\right)^{2n}}{n!}\underbrace{\left\{\int_{0}^{t}\exp\left(\tau-2t\right)\left[\int_{0}^{\tau}C_{\eta}(s)\mathrm{d}s\right]\mathrm{d}\tau\right\}^{n}}_{\tilde{C}^{n}(t)}$